### v2

# Multi-view Symbolic Regression

Abstract.
Symbolic regression (SR) searches for analytical expressions representing the relationship between a set of explanatory and response variables. Current SR methods assume a single dataset extracted from a single experiment. Nevertheless, frequently, the researcher is confronted with multiple sets of results obtained from experiments conducted with different set-ups. Traditional SR methods may fail to find the underlying expression since the parameters of each experiment can be different. In this work we present Multi-View Symbolic Regression (MvSR), which takes into account multiple datasets simultaneously, mimicking experimental environments, and outputs a general parametric solution. This approach fits the evaluated expression to each independent dataset and returns a parametric family of functions $f(x;\theta)$ simultaneously capable of accurately fitting all datasets. We demonstrate the effectiveness of MvSR  using data generated from known expressions, as well as real-world data from astronomy, chemistry and economy, for which an a priori analytical expression is not available. Results show that MvSR obtains the correct expression more frequently and is robust to hyperparameters change. In real-world data, it is able to grasp the group behaviour, recovering known expressions from the literature as well as promising alternatives, thus enabling the use SR to a large range of experimental scenarios.
genetic programming, symbolic regression, regression
copyright: noneccs: Computing methodologies Genetic programming
1. Introduction
The core of our modern scientific knowledge is based on careful production and analysis of experimental data. In a traditional scenario, the researcher’s task is to give meaning to measurable results, and analyze them in the light of an hypothesis. Frequently, the goal of this exercise is to find a mathematical description which can, at the same time, describe the recorded outcomes according to the state of the art of the field, and predict results from future similar experiments. Scientists became experts in the highly non-linear thought process required to interpret and translate scientific knowledge into suitable mathematical expressions. Due to the increase in data complexity, this task has also been approached through a machine learning perspective with the goal of automating the thought process.
Symbolic Regression (SR), for example, produces a mathematical expression with one or more variables that optimally fits a given data set. It searches for a parametric model $f(x;\theta)$ that minimizes a loss function measuring the goodness-of-fit to the input data. Primarily, it searches for the correct functional structure $f(x;.)$ and, within this process, also tries to establish the correct values of the parameters $\theta$. This method has been successfully applied to simulated scientific data sets in physics (Cranmer et al., 2020), chemistry (Hernandez et al., 2019), medicine (La Cava et al., 2023) and social sciences (de França et al., 2023), to cite a few. Traditionally, SR algorithms used ephemeral random constants to set these parameters values, but more recently, the algorithms apply a nonlinear optimization to fit the function to the data more accurately.
In many real scientific applications, the researcher is faced with different sources of data describing the same model but acquired from different viewpoints. In practice, this means that we have more data available to fit the model but, though they share the same functional structure, they may differ in parameter values. For example, consider data related to the contagion dynamics of a new virus collected from different populations. Even though the dynamics of the virus should be the same among all of the populations, the different aspects of a given population may influence the observed behavior in such a manner that all populations behave following a specific $f(x;\theta)$, but with different values of $\theta$. We argue that even if data of a single viewpoint (hereafter single-view) is abundant, multiple views can help constraining the search space of hypothesis.
In this paper, we introduce the idea of Multi-View Symbolic Regression (MvSR) that allows the practitioner to use the information from multiple sources and control the desired number of parameters to guide the search process to a parametric model that contains the correct number of parameters so it is neither too flexible (i.e., an universal approximator) nor too rigid (i.e., fitting only a single data source). This stimulates the discovery of scientific models that facilitates further analysis and interpretation of the phenomena. Specifically in this paper, we propose an adaptation to the Operon (Burlacu et al., 2020) package to calculate the fitness of a candidate solution by fitting this solution and calculating the base fitness individually on each data source and then returning an aggregated fitness of all sources. The experiments bring evidence that such procedure can successfully retrieve models capable of fitting all of the provided datasets.
Section 2 describes the algorithm and our adaptation to the fitness calculation. In Section 3 we describe the experimental methods used to assess the validity of our approach. Section 4 show the obtained results followed by a discussion. In Section 5 we apply MvSR to real scientific datasets from different fields and discuss the functional forms proposed. Finally, in Section 6 we conclude this paper with some final thoughts and future perspectives.
2. Multi-View Symbolic Regression
The goal of Multi-View Symbolic Regression (MvSR) is to search for a general parametric model that can simultaneously describe multiple data sets generated by the same underlying mechanism.
To illustrate this concept, imagine that we have data from multiple epidemic events from different countries occasioned by the same pandemic. If we aggregate these data into a single dataset and fit a regression model, we will obtain a model that returns the expected contamination for a given time $t$. This expected value is calculated over the sum of all events and can be seen as explaining the average contagion event worldwide. On the other hand, if we have a model $f(t;\theta)$ fitted independently for each event, the parameters values will reveal different aspects for each of them. For example, fitting this model to data from different countries will describe different reproduction rates and thus inform about the characteristics of local populations (i.e., widespread use of mask, density of the population, etc.).
In summary, MvSR finds a model $f(x;\theta)$ minimizing the aggregated error when fitting the parameters $\theta$ independently for each experiment, while applying a constraint on the number of parameters. More formally, given a dataset $\mathcal{D}=\left\{(x_{i},y_{i})\right\}_{i=1}^{p}$ with $p$ data points $x_{i}\in\mathcal{R}^{d},y_{i}\in\mathcal{R}$, SR seeks a parametric model $f(x;\theta)$ that minimizes a loss function $\mathcal{L}$ given the optimal parameter values $\theta$:
$\min_{f,\theta}\mathcal{L}(f(x;\theta),y).$
However with MvSR  the objective becomes:
(1) $\min_{f}{agg_{i=1..k}{\left(\min_{\theta^{i}}\mathcal{L}(f(x^{i};\theta^{i}),y^{i}\right)}},$
where the superscript $i$ refers to the index of each dataset, $agg$ is an aggregation function such as $max,avg,med$. Additionally, we impose the constraints
$x\in\mathbb{R}^{m}$
$\theta\in\mathbb{R}^{n}$
[…]
where $m$ is the number of independent variables ($x$), $n$ is the number of parameters in the model ($\theta$) with $n$ being bounded by a finite subset of the natural numbers.
Overall, given $k$ different datasets, we want to find the function $f$ with a limited number of parameters that minimizes the aggregated value of $\mathcal{L}$ for each dataset when independently adjusting the value of $\theta$ for each set. The purpose for these constraints is to find a model that can correctly adjust to the data without underfitting nor overfitting while containing the smallest number of parameters that summarizes each data set. A complete and ideal implementation of MvSR should allow to:
  1. (1)
    Control the maximum number of parameters.
  2. (2)
    Allow parameters to be repeated.
  3. (3)
    Receive multiple datasets as input.
  4. (4)
    Perform a minimization of the parameters independently for each dataset.
  5. (5)
    Use an aggregation function to compute an overall loss.
  6. (6)
    Penalise solutions based on the number of parameters used.
For this work, we present a proof of concept of the ideal MvSR approach where only the points 3, 4 and 5 of the list above have been fulfilled. Our current implementation (detailed in Section 2.1) slightly modifies a pre-existing Symbolic Regression algorithm at the evaluation phase. So, when a solution is evaluated, for each one of the $k$ datasets the algorithm will: fit the model to the datasets, calculate the losses, and then aggregate the results with the aggregation function. The algorithm will then use this result as a metric for reproduction (within the context of genetic programming).
2.1. Implementation details
In this work we use Operon (Burlacu et al., 2020), a high-performance C++ framework supporting single and multi-objective GP with non-linear optimization of the parameters using the Levenberg-Marquardt algorithm  (Kommenda et al., 2020). This framework was reported to perform well with respect to the runtime and overall quality of the results (La Cava et al., 2021). The framework also has a Python module counterpart, called PyOperon, which presents all of the flexibility of the C++ version.
The implementation was adapted to support the creation of independent evaluators, each of which evaluates the symbolic expression to one of the inputted datasets. Each evaluator fits the expression to one dataset, such that the expression can have different adjusted parameters for each view, evaluate the loss function (i.e., mean squared error, log-likelihood), and then aggregate the losses using an aggregation function. This final value is then used during selection and reproduction stages to calculate the probability of survival.
The available aggregation functions are average, median, min, max, harmonic mean. Notice that since Operon always minimizes the fitness, the min and max aggregation functions represent the best and worst fits, respectively. We use $max$ as the aggregate function, effectively making the objective as the function $f$ that minimizes the worst fit among the different sources of data. After the evolutionary process, the current implementation will return the symbolic expression fitted on the last dataset as a string. The best found expression is then converted into a parametric solution using SymPy (Meurer et al., 2017). The expression is simplified before replacing each float by a free parameter and converted into a python function. The particular float values of the last dataset are used as initial guess for future external minimization.
3. Experiments
To demonstrate the advantages of MvSR when multiple data sources are available, we devised an experimental design using data artificially created from the same generating function with either different parameters or covering different regions of its domain. Besides artificial benchmarks, we applied MvSR to three real-world datasets from different scientific fields and highlighted the benefit of using this approach instead of a traditional SR.
3.1. Data generation
For the artificial data, we setup a series of challenging benchmarks based on standards from the SR literature. For this purpose we chose three generating functions:
(2) $f_{1}(x)$ $=\theta_{0}+\theta_{1}x+\theta_{2}x^{2}+\theta_{3}x^{3}$
(3) $f_{2}(x)$ $=\sin{(\theta_{0}x_{0}x_{1})}+\theta_{1}(x_{2}-\theta_{2})^{2}+\theta_{3}x_{3}+x_{4}$
(4) $f_{3}(x)$ $=(\theta_{0}x_{0}^{2}+(\theta_{1}x_{1}x_{2}-\frac{\theta_{2}}{(\theta_{3}x_{1}x_{3}+1)})^{2})^{0.5}$
$f_{1}(x)$ is a third order polynomial function which constitutes a simple one dimensional case for which results can easily be visualised and interpreted. $f_{2}(x)$ and $f_{3}(x)$ are based on the Friedman functions (Friedman, 1991) for which free parameters have been added by replacing some constant values.
From the parametric functions, we generate examples which individually carries incomplete information about their parent function. For the polynomial function ($f_{1}$) we test two separate cases: i) each dataset is generated with two parameters set to 00 (see Views 1 to 4 in Table 1), ii) each dataset uses the same parameters (see partial view in Table 1) but displays only a narrow part of the behavior from which extrapolation of the original function is challenging. For $f_{2},f_{3}$ we will only test the first case where, by setting some the coefficients to 00, we mischaracterize the original function with the goal of rebuilding it using the different views.
Table 1. Parameter values used for each view and the partial view (only for $f_{1}$). For the different views we kept two parameters as 00 to depict the extreme situation where some parameters have no effect into the data. The partial has no parameters equal to 00 but each example is very restricted in the sampling range.
View 1 View 2 View 3 View 4 Partial view
$\theta_{0}$ $2$ 00 00 $2$ $2$
$\theta_{1}$ $2$ $2$ 00 00 $-2$
$\theta_{2}$ 00 $2$ $2$ 00 $2$
$\theta_{3}$ 00 00 $2$ $2$ $2$
Each of these benchmarks will contain a total of $4$ datasets (i.e., views) with the sample sizes fixed for each view. $f_{1}$ datasets consists of 20 points equally spaced on the interval $[-2,2]$. In the partial view case we keep the same number of points but we sample each dataset into evenly spaced domains $[-2,-1],[-1,0],[0,1]$ and $[1,2]$. $f_{2}$ and $f_{3}$ datasets consist of 100 points uniformly distributed on the intervals detailed in the scikit-learn package ${}^{,}$ [^2] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_friedman2.html.
To homogenize the results among the different benchmarks, we scaled the target variable $y$ of each dataset by applying the transformation $y^{\prime}_{i}=10\cdot\frac{y_{i}}{\max(\mathrm{abs}(y_{i}))}$. The factor 10 was arbitrarily chosen to multiply the absolute error, thus avoiding all scores to be clustered around 0 independently of the fit quality. This procedure introduces an extra scaling parameter for the $f_{2}$ and $f_{3}$ functions, increasing the true number of free parameters to $5$.
For each generative function, we create multiple datasets using different noise rates, $\left\{0.000,0.033,0.066,0.100\right\}$, in order to verify the robustness of this approach w.r.t. to noise. The noisy target is sampled from the distribution $\mathcal{N}(y,\sigma_{y}\sqrt{\frac{\tau}{1-\tau}})$, where $\tau$ is the noise rate.
3.2. Operon Hyperparameters and Post-processing
As mentioned in Section 2.1, we used an adapted version of pyOperon  [^3] https://github.com/heal-research/pyoperon/releases supporting the use of multiple datasets and aggregation function. For the following experiments we used the hyperparameters depicted in Table 2.
Additionally, we varied the hyperparameter max tree size from $5$ to $25$ with increments of $2$. This experiment will serve two purposes: i) having a baseline of models simpler than the original generating function since the functions require a minimum size of $7,11,14$ to be correctly represented; ii) test whether MvSR is prone to overfitting if given the freedom to expand the expression to larger sizes, thus providing opportunity to fit the noise term as well.
Each one of the $176$ combinations of benchmark functions ($f_{1}$ partial domains, and $f_{1},f_{2},f_{3}$ multiple views), noise level and maximum size represents a single instance of our set of benchmarks. For each instance we ran pyOperon main SR module for each one of the four views independently as well as with the MvSR adaptation. Each experiment was repeated with $100$ different random seeds. After each run, the string representation of the best symbolic model is processed with Sympy to replace the numerical values with parameter variables and the corresponding expressions are stored for post-processing.
These expressions are individually refitted to the noiseless version of each dataset minimizing the least squares with the python package iminuit (Dembinski and et al., 2020). The final score is calculated using the mean squared error (MSE). Therefore, if the correct expression (or any equivalent) is generated the score will be 00, even for functions generated on noisy datasets. When larger trees are allowed, some runs may generate NaN values as output. In such case the value is replaced by infinity.
Table 2. List of the fixed hyperparameters used in the experiments.
Parameter Value
population size
$1000$
number of evaluations
$100000000$
pool size
$5$
error metric
$MSE$
prob. cx
$1.0$
prob. mut.
$0.25$
max depth.
$10$
optim. iterations
$100$
aggregation function
max
operators
add, sub, mul, div, square, exp, sqrt, sin ($f_{2}$ only)
4. Artificial Benchmark Results
Figures 1 show heatmap plots of the obtained results for each benchmark function for every combination of noise level and maximum size. In these heatmaps, the colorscheme displays the average MSE of each combination, the lighter the color the better the result. Any values higher than 5 (including infinity) is depicted as the darkest color in order to keep the contrast in the visualization of the results. In each plot we show the worst and best single-view results (using the average of the plotted values as a choice criteria) and the MvSR results.
Figure 1. Heatmap of the average MSE of the tested combinations of noise and maximum expression size. Each row of heatmaps show results for the $f_{1}$, $f_{1}$ partial domains, $f_{2}$ and $f_{3}$ benchmarks, respectively. Columns represent the worst single-view (left), best single-view (center), and MvSR results (right). The colorbar represents the average MSE for that configuration, ranging from 00 (white) to a clipped value of $5$ dark blue. The clipping improves the comparison of small values.
From Figures 0(a), 0(b) and 0(c) we can see that for the noisy datasets and after a certain value of maximum size, both the single-view and MvSR behaves similarly reaching the correct solution. For the noiseless dataset, only MvSR is capable of finding the correct solution. This is expected since the single-view approach is submitted to a simplified version of the original expression and, thus, choose the simplest solution.
In Figures 0(d), 0(e) and 0(f) we can see a degradation of the results when using single-view with the noisy data. As the single-view has access to only one part of the polynomial curve, it can often fit the training data with a lower degree polynomial. This does not happen with the MvSR as it recovers a correct solution in almost every combination of noise and maximum size. The same behavior is displayed in Figures 0(g) to 0(l) where, in every combination, MvSR returns an optimal model (or near-optimal if the optimal model is not allowed by the maximum tree size) while the single-view approaches struggle with the lack of information and higher noise levels. We observe that MvSR is resilient against noise and successfully prevents overfitting even when the maximum size is larger than the original expression.
Another important question is whether MvSR returns a model with the correct number of parameters. A model with less than the optimal number of parameters is less flexible to fit the data from different sources and thus, has a tendency to underfit. Similarly, a model with more parameters than needed, can be too flexible and capable of fitting the noise imbued in the data, thus creating an overfitted model.
Having the correct number of parameters is important when building a model to not only avoid under and overfitting, but also enable model interpretation. Figure 2 shows the critical difference diagram of the average rank calculated by comparing the absolute value between the optimal number of parameters and the number of parameters of each model. If the absolute difference is 00 it means the model contains the exact number of parameters, anything higher than that means that it either has more or less than the reference number. This plot shows that MvSR average rank (approx. $2$) is better than the single-view approaches. While the single-view approaches are indiscernible between each other.
Figure 2. Critical difference diagram of the average rank w.r.t. the absolute difference between the number of parameters of a benchmark and the number of parameters of the model. This diagram was built calculating the Friedman hypothesis test with $\alpha=0.05$ and Holm–Bonferroni correction. The names ex1, ex2, ex3, ex4 refers to the four single-view models.
5. Scientific application
In this section we apply MvSR to real experimental datasets coming from 3 scientific fields, namely chemistry, finance and astrophysics. Such problems represent significantly harder challenge for SR algorithms as the data generally have no absolute “correct” generative model, is irregularly sampled and display non-Gaussian noise. However it constitute perfect testing grounds for MvSR, since one cannot know a priori if examples provide display the full or partial behavior of the best parametric model. Therefore, in such cases, the most conservative approach is to always use MvSR in order to build general laws.
5.1. Chemistry dataset
Beer’s law is an empirical law widely used in chemistry which relates the attenuation of light to the properties of the material through which the light is travelling. In chemistry, the attenuation of a beam of light going through a solution is presumed to be only due to absorption, as solutions do not scatter light of wavelengths frequently used in analytical spectroscopy. In UV-visible spectroscopy, we characterize a solution by its transmittance. This corresponds to the ratio of light intensities before and after passing through the sample. We express the absorption of a solution as $A=-\log(T)$. We experimentally observe that all molecules display a common pattern: for $A\leq 1$, the absorption rises linearly with the concentration, while for $A>1$ the linearity breaks and the absorption increases more and more slowly until it reaches a plateau due to the limit of measurements of the spectrophotometer.
The Beer’s law is used to describe the properties of chemical species when $A\leq 1$. It states that the absorptive capacity of a dissolved substance is directly proportional to its concentration in a solution and is expressed as $A=\epsilon lc$ where A is the absorbance, $\epsilon$ is the molar extinction coefficient, l is the optical path length and c the concentration. A handful of alternative models have been proposed. They attempt to build a general Beer’s law capable of fitting a larger range of absorption. For example Bozdoğan (2022) suggested to use a quadratic polynomial equation to compensate the positive or negative deviations from the linearity. Yeh et al. (2023) proposed to extend law by adding two exponent parameters on $l$ and $c$, thus adding flexibility to describe the deviations. Both approaches extend the range of validity of the law but does not provide a solution general enough to properly characterize the absorption at any A value.
We propose to use MvSR on a set of measurements to find such a general parametric solution. In order to proceed, several wavelength scans were carried out using the Hitachi double-beam spectrophotometer UH3500 from 800 nm to 200 nm. Thus, four molecules were analyzed at various concentrations in dichloromethane, a commercial coumarin, two bodipy (Song et al., 2012)(Tran et al., 2023) and a porphyrin (tpp, 1967) which was previously synthesized in order to measure their absorbance A as a function of the solution’s concentration. Ideally the parametric model should recover their extinction coefficient $\epsilon$.
After a small hyper parameter exploration we produce a simple accurate parametric function using MvSR. It was produced with a max tree length of size 15 and the $exp$ and $log$ operators allowed.
(5) $f(x;\mu,\epsilon)=\log\left(\frac{1}{\mu+\mathrm{e}^{-\epsilon x}}\right)$
The data acquired display high non-linearity behaviors, in particular the absorption tend to reach a plateau around $A=3$. Figure 3 presents the best fits along with the parameter values associated and the $R^{2}$ score. The extended Beer’s law proposed by MvSR carries ideal properties to fit the observations. Indeed, in the case where the linearity is respected, the $\mu$ parameter can be set to 0 and the equation simplifies into the original Beer’s law. In such case $\epsilon$ carry exactly the same information as the standard method. In the general case, $1/\mu$ characterizes the plateau that the absorption will reach at high concentrations. MvSR shows that an exponential transition between the linear evolution and the plateau provides an accurate fit to the data. We observe that all $R^{2}$ scores are close to 1, demonstrating that the law is fitted to the observations. Additionally, Figure 3 displays the classical Beer’s law fitted to the data points for which $A\leq 1$. We also observe that the lower the extinction coefficient is, the further apart the two models are at low absorption. This suggests a deviation from Beer-Lambert’s law at low $\epsilon$, and would require further investigation.
Figure 3. Best MvSR fit (Equation 5) of the absorption as a function of the molar concentration for 4 different molecules. Gray lines correspond to the Beer’s law fitted to the data points for which $A\leq 1$.
In summary Equation 5 offers the possibility of computing molar extinction coefficients without being strictly restricted to the linear regime. The functional form proposed by MvSR contains two parameters and thus requires only a few data points to be constrained. It could be an easiest and quickest alternative for the determination of intrinsic properties of chemical species.
5.2. Finance dataset
Financial markets exhibit complex emergent behaviours. We define the return as the difference between the price of an asset at time $t$ and time $t+1$ is $r_{t}=p_{t+1}-p_{t}$. At first approximation, we can describe the distribution of returns by a Brownian motion (Bachelier, 1900). The field of econophysics, which applies tools from statistical physics to study these stochastic processes, seeks to identify probability distribution functions that go beyond such Gaussian approximation. Those initial models led to the development of the renowned Black-Scholes equation (Black and Scholes, 1973), but neglected the significance of rare and extreme events due to their lack of fat tails. Indeed, the distribution of these specific events can be fitted with power-laws (Mantegna and Stanley, 1995).
Modern models, as presented in (Mantegna and Stanley, 1999), focus on Lévy processes. These include, for example, the Gaussian distribution as well as the Cauchy distribution, the second being an already improved solution with its fat tails. Such models were proposed after studying the statistical properties that most data exhibit, such as the famous $S\&P500$ dataset, and then finding distributions that possess such properties. However, our Multi-view approach enables us to identify a common distribution for all assets by considering each of them individually. In this Section, we show that MvSR rediscover some of the presented distributions and propose new models that better fit the real return distributions.
We analyse time series data consisting of multiple prices generated by various assets. Each asset’s price corresponds to the value of a share of a company on financial markets. We utilise a publicly available Kaggle dataset  [^4] https://www.kaggle.com/datasets/iveeaten3223times/massive-yahoo-finance-dataset containing data from $491$ companies. The first 10 companies are used as views for MvSR, while the remaining companies are only used for testing purposes. Values of assets are taken each day at open market time, over a period of 5 years starting from January $1^{st}$, $2018$. For each asset, we study the distribution of its returns with a sampling of $100$ bins of equal width. We normalize the data as described in Section 3.1. This type of data exhibits common statistical properties. A positive mean that corresponds to overall economic growth. A leptokurtic profile that indicates more extreme events than a normal distribution would produce, and a negative skewness due to an asymmetric shape that informs rare events are more likely to be crises than economic booms. The first and last properties are only observed for datasets covering long time periods (months or years) such as ours.
Under those conditions, we explore multiple seeds, tree lengths (ranging from size 8 to 20) and operators (using various combinations of $exp$, $abs$ and $power$) to obtain multiple possible parametric solutions with MvSR. In Table 3 we present 6 parametric models generated by MvSR. We measure the performance of all the functions by fitting them to each asset and display the median MSE value as a comparative metric in the last two columns of Table 3.
In our experiment, MvSR noticeably recovers 3 widely used parametric forms: the Gaussian, Cauchy and Laplace distributions. In addition to these results, we obtain solutions similar to Laplace with some variations. These three distributions are presented in Table 3. We can see from their scores that those parametric forms present better fits than the solutions already present in the literature.
Figure 4. Normalized (Section 3.1) distribution of returns for 2 assets, fitted by the Cauchy model and the best MvSR solution (Power-Laplace).
Figure 4 shows the distributions of returns of some assets alongside the fit given by the best distributions, new and already found in the literature, that we obtain. We call the latter a Power-Laplace distribution  [^5] That we name after it’s functional composition $h(x)=g(f(x))$ with the Laplace distribution $g(x)=e^{-a\times|x|}$ and a power function $f(x)=x^{b}$.. It performs well because the power is always a root  [^6] We always find $10<b<$ 1., which accentuates the fat tail effect of the distribution, allowing a better fit in the tails while maintaining a good peaked fit in the centre.
Models Equation f(x) $med(MSE)$ $MSE_{S\&P}$
Gaussian (Bachelier, 1900; Black and Scholes, 1973) $A\cdot\mathrm{e}^{-\frac{x^{2}}{B}}$ 0.363 0.260
Laplace (Kou, 2002) $A\cdot\mathrm{e}^{-B|x|}$ 0.342 0.084
Cauchy (Liu et al., 2012) $A\cdot B^{2}/(x^{2}+B^{2})$ 0.305 0.079
Linear-Laplace $(A-Bx)\cdot\mathrm{e}^{-C|x|}$ 0.327 0.065
Exp-Laplace $A\cdot\mathrm{e}^{Bx-C|x|}$ 0.328 0.063
Power-Laplace $A\cdot\mathrm{e}^{B|x|^{C}}$ 0.246 0.075
Table 3. Best functions generated by MvSR. The last two columns respectively show the median MSE score of the functions fitted on individual normalized assets and the score when fitted on normalized the $S\&P500$ dataset. Bold numbers correspond to the best score of the column.
The last column of Table 3 shows the scores of all the functions fitted on the $S\&P500$ index. This index is an aggregated measure of the value of the assets of the 500 largest American companies. As such, the distribution of its returns should behave like an average distribution of these assets. For these data, the best distribution is the one in the second to last line of Table 3. Like for the previous application on multiple assets, this shows our new distributions are also better fits than those present in the literature.
Finally, we see that distributions found in the literature are relevant for the $S\&P500$, as they all perform much better than the Gaussian. However, apart from the Gaussian, their scores do not differ much. This shows that they are almost all equivalent fits for this aggregated index. On the other hand, their median scores for a variety of assets are much more disparate. In this second case, our algorithm produced much more efficient distributions ; distributions that outperformed all the others. This indicates that our multi-view approach is an efficient strategy for finding a unique distribution that gives good fits for as many assets as possible. We illustrate this effect on Figure 4. The right panel shows how Cauchy and Power-Laplace both fit well the $S\&P500$ index. The left panel displays a special case selected for its high score difference between both models. This shows how a model like Cauchy fails to fit some distributions compared to the best model found by MvSR, which is more versatile.
We showed that applying MvSR to a set of assets all at once, rather than looking at them separately or using an aggregated index, is an efficient approach to gaining new insights into markets behaviour and improving characterisation of the stochastic process that governs them. The results presented here could be fine tuned by adding more functions to the MvSR. For example the Gamma function could allow us to generate more distributions such as the Gamma variance process (Madan et al., 1998) and Student’s t-distribution (Bouchaud and Potters, 2000), which are the general forms for the distributions we already found using MvSR. Finally, further investigation of the distributions found here may prove insightful for future modelling in the field.
5.3. Astrophysics dataset
Astrophysics constitutes an ideal testing environment for MvSR as it encompasses a diverse array of non-linear equations aimed at modeling observations of large-scale phenomenon within the universe. Classical SR has been widely used to characterise various relationships such as photometric redshift(Krone-Martins et al., 2014), the Hertzsprung–Russell diagram or the plane of elliptical galaxies (Graham et al., 2013). After this project was completed, we also became aware of (Tenachi et al., 2023), who approached a similar problem when dealing with the treatment of stellar streams.
Astrophysical phenomena present a diverse landscape due to the vast range of physical values in stellar objects and the variety of physical processes occurring within them. One method to study these phenomena involves tracking the change in radiation flux over time. This time-series of flux measurements, known as a light curve, typically varies in the number of observations depending on the object and the optical filter in use. This is particularly true for ground-based surveys, such as the Zwicky Transient Facility (ZTF, (Bellm et al., 2019)), because of weather conditions, seasons and other factors.
The largest catalogs of variability classifications contain millions of stellar objects (Watson et al., 2006; Eyer et al., 2023), providing researchers an access to numerous light curves of the same type of variability. One of the most studied types of extreme variability is supernovae (SNe), a catastrophic cosmic event caused by a star’s explosion. The physical processes in a supernova are highly complex, and we may not yet be able to directly model individual supernova light curves. This makes such astrophysical data an excellent case study for MvSR.
We choose three Supernova type Ia (SNIa) with a good time sampling from ZTF Data Release 17: SN2019fck, SN2018aye and SN2021mwb (Bellm et al., 2019; Malanchev et al., 2023). We use observations in $g$ and $r$ photometric filters independently, resulting in a dataset of six examples for MvSR. The light curve shape of SNIa in $g$ filter could be roughly described as an exponentially rising brightness followed by an exponential fading. Such behavior is typically well modeled by parametric equations found in the literature (Bazin et al., 2009; Villar et al., 2019; Sánchez-Sáez et al., 2021; Russeil et al., 2023). However, being observed in a redder $r$ filter, SNIa light curves have a secondary bump. This particular behaviour is much harder to describe with a simple parametric equation. We choose to use both $g$ and $r$ band to provide a wide variety of examples to MvSR.
We apply data quality cuts and select only data points with a signal to noise ratio greater than 20. The data was shifted so that the observed peak time is at 0 and the entire light curve was normalized by its maximum flux. Finally we consider only data points with a time ranging from -50 to 150 days from the observed peak.
Equation f(t) $<R^{2}>$ No. parameters
$\mathrm{e}^{-At\cdot(B-\mathrm{e}^{-Ct})}$ 0.990 3
$\frac{A}{(B\cdot\mathrm{e}^{Ct}+\mathrm{e}^{-Dt})}$ 0.987 4
$\frac{A^{Bt}}{Ct+(-Dt+\mathrm{e}^{Et})^{2}}$ 0.992 5
Table 4. Summary of the best parametric functions generated using MvSR on SNIa lightcurves. The second column corresponds to the mean $R^{2}$ score over the 6 examples provided.
We use a maximum tree length of 12, include the $exp$, $square$ and $power$ operators and explore multiple seeds to obtain a panel of possible parametric solutions. Table 4 displays a selection of the best parametric solutions found by MvSR along with the mean $R^{2}$ score of the best fit on the SNe.
Noticeably, the second equation is already known as the  Bazin function (Bazin et al., 2009) and is widely used in the literature (Russeil et al., 2023)(Dai et al., 2018). However, the standard form includes a $t_{0}$ parameter appearing twice which is added to t, effectively encoding for a time shift. The form generated by MvSR is mathematically equivalent but uses parameters appearing only once. The first solution presented in Table 4 is characterized by an intricate exponential form. Despite lacking similar counterparts in the literature, it provides excellent fits and even outperforms the standard Bazin function (Figure 5). Moreover, since it uses only 3 parameters, additional ones such as $t_{0}$ or a scaling parameter could be added to make it more versatile. Finally the last equation in Table 4 requires 5 parameters and provides the most accurate description of the SNe. MvSR can produce solutions of arbitrary size with increasingly good fit, however we choose to limit it to 5 parameters maximum in order to prevent the overfitting.
Figure 5. Best fit of two parametric functions found by MvSR on SN2021mwb in the $g$ and $r$ filters. The right panel corresponds to the Bazin function commonly used in the literature.
Overall MvSR was able to generate multiple good models to describe SNe behaviors. It recovered a solution from the literature and even proposed improved solutions in terms of goodness of fit and/or number of parameters used. However, the generated models struggle with the same problem as the equations from the literature: they don’t provide a simple description of the second bump of the SNIa in the $r$ band. It may also highlight a limitation of the current MvSR implementation, which doesn’t include parameter repetition, as stated in Section 2). Given that a repeated time shift parameter is standard in models used in the literature, this results highlights the importance of a more complete MvSR implementation for practical scientific applications.
The code used to produce all the results presented in this work is publicly available [^7] https://github.com/erusseil/MvSR-analysis.
6. Conclusions
Symbolic Regression (SR) has proven to be extremely efficient in searching for mathematical expressions that describe the relationship between a set of explanatory and response variables. In its traditional form, it translates the behavior of one such data set into an analytical function which can be used for further analysis. Nevertheless, in a realistic scenario, the researcher is frequently faced with multiple outcomes from the same experiment. These may correspond to different experimental setups, initial conditions or domain coverage, but are all generated by the same underlying mechanism, one which the researcher aims to describe.
In this work, we proposed Multi-View Symbolic Regression (MvSR), a framework that exploits this scenario by extending the scope of traditional SR, allowing the user to provide multiple examples as input. The algorithm searches for the best parametric function which simultaneously describes all the input data provided. This is achieved by fitting each input data set independently with the same regression model and aggregating their individual fitness into a single one.
For this purpose, we included an aggregation fitness function in Operon that supports different aggregators such as mean, maximum, median, or harmonic mean.
We have tested this approach with four different challenging artificial benchmarks with added noise and composed of extreme situations where some of the parameters are set to 00 or the domain coverage is limited. When compared to the traditional approach we report that MvSR is capable of correctly retrieving the original expression in most configurations with a higher accuracy than SR, even in a presence of a strong noise. Additionally, we stress-tested our method by using real-world experiments from three different areas: chemistry, finance and astrophysics. MvSR was not only capable of recovering well known models from the literature but it also found new alternatives that are promising in these fields.
Our results showcase the potential enclosed in applying MvSR in real scientific scenarios. These could be made even better with further functionalities, like enabling a maximum number of parameters to be used in the model, either as a hard constraint or through a penalization term, and allowing the same parameter to appear more than once in the final expression. Such additional features would not only result in more flexible functions, with smaller number of parameters, but it would also allow the researcher to indirectly tailor the final result, thus increasing the chances of a parametric form which can inspire interpretability.
In this context, a full implementation of MvSR would have much broader applications than the ones described here. We intend to further explore its ramifications in dedicated studies focused on the particularities of each science case. Nevertheless, in order to fully exploit its potential and popularize its use within a broad range of scientific areas, a user-friendly implementation of all the above mentioned functionalities is paramount.
Acknowledgements
This project emerged from discussions during two SNAD workshops: SNAD-IV [^8] https://snad.space/2021/ and SNAD V [^9] https://snad.space/2022/, held in France, Briançon and Clermont Ferrand, respectively. ER, KM and EEOI thank the entire SNAD team for the insightful discussions and environment. ER acknowledges the ISITE grant from the ”AAP Séjour-recherche” program from Université Clermont-Auvergne (UCA). FOF acknowledges support from Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), grant number 2021/12706-1 and CNPq through the grant 301596/2022-0. EEOI thanks Rafael S. de Souza and Alberto Krone-Martins for the introduction to Symbolic Regression. Finally we thank the pastéis for being a true source of inspiration.